A sample of 35 observations is selected from a normal population. The sample mean is 20, and the population standard deviation is 2. Conduct the following test of hypothesis using the 0.05 significance level. H0: μ ≤ 19 H1: μ > 19

  1. What is the value of the test statistic? (Round your answer to 2 decimal places.)

What is the p-value? (Round your answer to 4 decimal places.)

  1. e-2. Interpret the p-value? (Round your final answer to 2 decimal places.)
Answer #1

Part c)

Test Statistic :-
Z = ( X - µ ) / ( σ / √(n))
Z = ( 20 - 19 ) / ( 2 / √( 35 ))
Z = 2.96

P value = P ( Z < 2.958 ) = 0.0015

Looking for the value Z = 2.96 in standard normal table to find the P value.

Interpretation

A small p-value indicates strong evidence against the null hypothesis, so you reject the null hypothesis.

A large p-value indicates weak evidence against the null hypothesis, so you fail to reject the null hypothesis.